OpenStudy (anonymous):

State the amplitude, period, phase shift, and verticle translation for y = -sin(x-pi/4)+2

OpenStudy (anonymous):

\[y = -\sin(x-\frac{ \pi }{ 4 })+2\]

OpenStudy (anonymous):

@TeresaCal Last trig function problem, thanks a ton for the help on the last one :)

hero (hero):

The general formulas for sine and cosine are y = Asin(Bx C) + D and y = Acos(Bx C) + D. We shall begin by labeling the various components. A amplitude of the function = (max-min)/2 (vertical stretch of the graph) B stretch/shrink on the x-axis. (It compresses or expands the graph.) C/B the phase shift of the graph (the shift left (if C/B is neg.) or right (if C/B is pos.)) D the vertical shift of the graph.

OpenStudy (anonymous):

Starting with amplitude, A = 1 correct?

hero (hero):

Anything before "sin" is included with A. So A = -1

OpenStudy (anonymous):

Alright, now B. B = 1 right?

OpenStudy (anonymous):

I would assume 1 just takes the place considering there is nothing there. Sorry, I'm not used to seeing functions set up like this.

hero (hero):

Yes, B = 1 in this case

OpenStudy (anonymous):

Okay, and I know to find the period it is 2pi/B so in this case, the period is simply 2pi correct?

hero (hero):

Correct again

OpenStudy (anonymous):

Excellent, now for the phase shift. Would the answer for phase shift be written simply as "pi/4 units to the right"?

hero (hero):

Sounds about right :)

OpenStudy (anonymous):

Alrighty and vert shift is 2 units up of course

hero (hero):

Hang on...

OpenStudy (anonymous):

Oh wait, phase shift would be horizontal compression right?

hero (hero):

No, don't get those two things confused.

hero (hero):

Phase Shift is C/B In this case, B = 1 so C/B = pi/4

hero (hero):

If B was anything other than 1, then the phase shift would be different from the horizontal compression

hero (hero):

Only when B = 1 does phase shift = horizontal compression

OpenStudy (anonymous):

Okay okay :) One of the trickiest parts for me is keeping the whole horizontal/verticle compression and stuff.

OpenStudy (anonymous):

Well that just about does it @Hero Thank you so much for all the help and I wish I could give more medals!

hero (hero):

One medal is good enough, thanks :)